OpenStudy (anonymous):

What is the value of the expression when a = 2? 5a^4

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

replace "a" with 2 what is a^4?

OpenStudy (anonymous):

8

Vocaloid (vocaloid):

nope, remember what I said before...

Vocaloid (vocaloid):

^ is an exponent...

OpenStudy (anonymous):

2*2*2*2

Vocaloid (vocaloid):

right, and that equals?

OpenStudy (anonymous):

16. My bad.

OpenStudy (anonymous):

Someone told me it was 10000. It's not though.

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

good, so if a^4 = 16, what is 5(a^4)?

Vocaloid (vocaloid):

and sorry for the delays, this website lags like crazy

OpenStudy (anonymous):

So 5*5*5*5=625

OpenStudy (anonymous):

And your fine. So 625 is my answer.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Or is there more I need to do?

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@mathstudent55

Vocaloid (vocaloid):

uh not quite

OpenStudy (anonymous):

Alright. What d I need to do next?

Vocaloid (vocaloid):

a^4 = 16 5(a^4) = 5(16) = ?

OpenStudy (anonymous):

5(a^4)= 625 5(16)=80

OpenStudy (anonymous):

625+80=705

OpenStudy (anonymous):

is 705 my answer?

Vocaloid (vocaloid):

no it's just 80